# SG-I2V: Self-Guided Trajectory Control in Image-to-Video Generation

### v2

@@ 3.3 Trajectory Control with Latent Optimization
where $\odot$ is Hadamard product, and $\tilde{\bm{F}}_{n}(\bm{z}_{t})[\mathcal{B}_{b,n}]\in\mathbb{R}^{h_{b}\times w_{b}\times d}$ is feature maps cropped by the bounding box $\mathcal{B}_{b,n}$. Following DragAnything (Wu et al., 2024c), we weigh the feature difference using a Gaussian heatmap $\bm{G}_{b}\in\mathbb{R}^{h\times w}$. This focuses on optimizing pixels closer to the bounding box center, as pixels near the edge may be background pixels that we do not want to move.
@@ 4.2 Qualitative Results
Fig. 1 presents the versatile control ability of our method. We can control foreground objects to perform rigid motions, such as trains moving, and non-rigid motions, such as the movement of human hairs. In addition, we can control non-physical entities such as smoke and fire. The moved scene elements naturally adapt to the new location while preserving their original identity. Thanks to our general formulation of trajectories, camera motion control is also supported. We highly encourage the readers to view our project page for additional results.
Figure 4: Failure cases in supervised baselines. We observe that DragNUWA tends to distort objects rather than move them, and DragAnything is weak at part-level control as it is designed for entity-level control. In contrast, our method can generate videos with natural motion for diverse object and camera trajectories. Please see our project page for more general comparisons.
@@ 4.4 Ablation Studies
Fig. 10 summarizes the effects of optimizing features across different denoising timesteps ($t=50$ corresponds to standard Gaussian noise). Lower timesteps (e.g., $t=20,1$) significantly impair visual quality, motivating us to optimize only on earlier denoising steps. Conversely, timesteps greater than $t=45$ degrade motion fidelity due to the lack of detailed semantic information at extremely high noise levels. Notably, ObjMC improves at $t\in[20,1]$, but due to the Co-tracker tracking moving artifacts (See here). Based on these observations, we optimize feature maps extracted between timesteps 30 and 45, balancing visual quality and motion fidelity.
@@ 5 Conclusion
Limitation and future works.
@@
Currently, our synthesized videos are not yet indistinguishable from real footage. Still, we anticipate that future iterations of large video models will achieve even greater quality, potentially producing videos that look photoreal. This prospect brings significant ethical and safety considerations, as cutting-edge generative models can be misused to ill effect. We advocate against misuse of generative models to malign or misinform, and we encourage safe and responsible use of such models Google (2023).
@@ Appendix A Experimental Setup Details
Details of FVD. As mentioned in Sec. 4.1, we resize all the generated videos to $320\times 576$ before inputting into the evaluation scripts to ensure a fair comparison. During the computation of FVD, we further resize all the videos to $256\times 256$ following DragAnything (Wu et al., 2024c).
@@
Details of ObjMC. Following DragAnything (Wu et al., 2024c), we compute ObjMC to evaluate motion fidelity. ObjMC is defined as the framewise average distance between the generated and ground truth trajectories, where the ground truth trajectories are reused from DragAnything and the generated trajectories are estimated using Co-Tracker v2 (Karaev et al., 2024). Since the VIPSeg dataset comprises videos with varying resolutions  (e.g., $480\times 800$, $1440\times 2560$), we resize all the videos to a uniform resolution (as specified in Tab. 1) before inputting to the diffusion model. Unlike DragAnything, which resizes the generated videos back to the original resolutions to compute ObjMC, we resize all of them to a fixed resolution of $320\times 576$. Additionally, we exclude ground truth trajectory points that are not within the image space due to objects moving out of frame. Furthermore, we exclude short videos with fewer than $14$ frames for evaluation.
@@
Details of zero-shot adopted methods. Due to the unavailability of zero-shot trajectory-controlled image-to-video generation, we adopt techniques from text-to-image or text-to-video methods to create new baselines in Tab. 1.
@@
FreeTraj incorporates the noise initialization technique known as trajectory injection from (Qiu et al., 2024). This work has observed that introducing motion bias into the initial noise, by copy-pasting the noise from the first frame to other frames along the trajectories, influences the motions in the generated videos of text-to-video diffusion models. However, we did not observe similar effects in image-to-video diffusion models; instead, we observed a degradation in visual quality, likely due to the latents being out of distribution.
@@
DragDiffusion is inspired by the image editing technique in (Shi et al., 2024), which utilizes feature maps extracted from the output of upsampling blocks to guide the generation process. Concretely, we perform optimization on the feature maps output from the second and third upsampling blocks located at the middle resolution level. Although the original DragDiffusion does not include high-frequency preserved post-processing, we incorporate this in the baseline to enhance visual quality. We use the same hyperparameters as in our method.
@@
MOFT is similar to DragDiffusion but optimizes on feature maps derived from Content Correlation Removal proposed by Xiao et al. (2024a). Concretely, we optimize on the outputs of the upsampling layers subtracted by their mean features averaged across frames. These feature maps are well-suited for reference-guided motion generation as they are motion-consistent (i.e. pixels with similar motion exhibit similar feature vectors) enabling direct comparison with reference motion feature maps. However, these feature maps are less sensitive to semantic information, and therefore, are not effective for our optimization, where reference motion is unavailable.

### v3

@@ 3.3 Trajectory Control with Latent Optimization
where $\odot$ is Hadamard product, and $\tilde{\bm{F}}_{n}(\bm{z}_{t})[\mathcal{B}_{b,n}]\in\mathbb{R}^{h_{b}\times w_{b}\times d}$ is feature maps cropped by the bounding box $\mathcal{B}_{b,n}$. Following DragAnything (Wu et al., 2024c), we weight the feature difference using a Gaussian heatmap $\bm{G}_{b}\in\mathbb{R}^{h\times w}$. This focuses on optimizing pixels closer to the bounding box center, as pixels near the edge may be background pixels that we do not want to move.
@@ 4.2 Qualitative Results
Fig. 1 presents the versatile control ability of our method. We can control foreground objects to perform rigid motions, such as trains moving, and non-rigid motions, such as the movement of human hairs. In addition, we can control non-physical entities such as smoke and fire. The moved scene elements naturally adapt to the new location while preserving their original identity. Thanks to our general formulation of trajectories, camera motion control is also supported. We highly encourage the reader to view our project page for additional results.
Figure 4: Failure cases in supervised baselines. We observe that DragNUWA tends to distort objects rather than move them, and DragAnything is weak at part-level control as it is designed for entity-level control. In contrast, our method can generate videos with natural motion for diverse object and camera trajectories. Please see our project page for additional comparisons.
@@ 4.4 Ablation Studies
Fig. 10 summarizes the effects of optimizing features across different denoising timesteps ($t=50$ corresponds to standard Gaussian noise). Lower timesteps (e.g., $t=20,1$) significantly impair visual quality, motivating us to optimize only on earlier denoising steps. Conversely, timesteps for $t>45$ degrade motion fidelity due to the lack of detailed semantic information at extremely high noise levels. Notably, ObjMC improves at $t\in[20,1]$, but due to the Co-tracker tracking moving artifacts (e.g., see project webpage). Based on these observations, we optimize feature maps extracted between timesteps 30–45 to balance visual quality and motion fidelity.
@@ 5 Conclusion
Limitations and future work.
@@
Acknowledgements
DBL acknowledges support from NSERC under the RGPIN and Alliance programs, the Canada Foundation for Innovation, and the Ontario Research Fund.
@@ Acknowledgements
Currently, our synthesized videos are not yet indistinguishable from real footage. Still, we anticipate that future iterations of large video models will achieve even greater quality, potentially producing videos that look photoreal. This prospect brings significant ethical and safety considerations, as cutting-edge generative models can be misused to ill effect. We advocate against misuse of generative models to malign or misinform, and we encourage safe and responsible use of such models (Google, 2023).
@@ Appendix A Experimental Setup Details
Details of FVD. As mentioned in Sec. 4.1, we resize generated videos to $320\times 576$ before feeding them into the evaluation scripts to ensure a fair comparison. For FVD computation, we further resize all videos to $256\times 256$, following DragAnything (Wu et al., 2024c).
@@
Details of ObjMC. Following DragAnything (Wu et al., 2024c), we compute ObjMC to evaluate motion fidelity. ObjMC is defined as the framewise average distance between the generated and ground truth trajectories, where the ground truth trajectories are reused from DragAnything, and the generated trajectories are estimated using Co-Tracker v2 (Karaev et al., 2024). Since the VIPSeg dataset contains videos with varying resolutions (e.g., $480\times 800$, $1440\times 2560$), we resize all videos to a uniform resolution (as specified in Tab. 1) before feeding them into the models.
Unlike DragAnything, which resizes generated videos back to their original resolutions for ObjMC computation, we resize all videos to a fixed resolution of $320\times 576$. Additionally, we exclude ground truth trajectory points that fall outside the image space due to objects moving out of frame. We also omit short videos with fewer than 14 frames from the evaluation.
@@
Details of zero-shot adopted methods. Due to the unavailability of zero-shot trajectory-controlled image-to-video generation methods, we adopt techniques from text-to-image and text-to-video methods to establish new baselines in Tab. 1.
@@
FreeTraj incorporates the noise initialization technique (called trajectory injection) from (Qiu et al., 2024). This work has shown that introducing motion bias into the initial noise—by copy-pasting the noise from the first frame to other frames along the trajectory—can influence motion in the generated videos of text-to-video diffusion models. However, we did not observe similar effects in image-to-video diffusion models. Instead, we found that this approach degraded visual quality, likely because the modified input latents fell out of distribution.
@@
DragDiffusion is inspired by the image editing technique in (Shi et al., 2024), which utilizes feature maps extracted from the outputs of upsampling blocks to guide the generation process. Specifically, we optimize on the feature maps output by the second and third upsampling blocks at the middle resolution level. Although the original DragDiffusion does not include high-frequency preserved post-processing, we incorporate it in our baseline to enhance visual quality. We use the same hyperparameters as in our method.
@@
MOFT is similar to DragDiffusion but optimizes feature maps derived from Content Correlation Removal, as proposed by Xiao et al. (2024a). Specifically, we optimize on the outputs of the upsampling layers subtracted by their mean features averaged across frames. These feature maps are well-suited for reference-guided motion generation because they are motion-consistent (i.e., pixels with similar motion exhibit similar feature vectors), enabling direct comparison with reference motion feature maps. However, they are less sensitive to semantic information and, therefore, ineffective for our optimization, where reference motion is unavailable.